OpenStudy (destinykiara99):

Find a cubic function with the given zeros. Square root of seven., -Square root of seven., -4 f(x) = x^3 + 4x^2+ 7x - 28 f(x) = x^3 + 4x^2 - 7x - 28 f(x) = x^3 - 4x^2 - 7x - 28 f(x) = x^3 + 4x^2 - 7x + 28

OpenStudy (sshayer):

\[f(x)=\left( x-\sqrt{7} \right)\left( x+\sqrt{7} \right)\left( x+4 \right)\] simplify it.

OpenStudy (destinykiara99):

@ganeshie8 @jim_thompson5910 @sweetburger @mathstudent55 @sammixboo

OpenStudy (destinykiara99):

@hartnn

OpenStudy (destinykiara99):

PLEASE HELP, I AM RUNNING OUT OF TIME TO FINISH

OpenStudy (lilah):

@jim_thompson5910

OpenStudy (lilah):

@ganeshie8

jimthompson5910 (jim_thompson5910):

as @sshayer wrote out, expand that polynomial out to get... \[\Large f(x) = (x-\sqrt{7})(x+\sqrt{7})(x+4)\] \[\Large f(x) = (x^2-(\sqrt{7})^2)(x+4)\] \[\Large f(x) = (x^2-7)(x+4)\] \[\Large f(x) = x^2(x+4)-7(x+4)\] I'll let you finish up

OpenStudy (destinykiara99):

f(x)=x^3+4x^2-7x-28 ???

OpenStudy (lilah):

yes that's right

OpenStudy (lilah):

that's the answer

jimthompson5910 (jim_thompson5910):

I agree

OpenStudy (lilah):

ik because i used mathway, an online calculator u should check it out it helps a lot

OpenStudy (destinykiara99):

Thanks :)